# Team Xiaomi EV-AD VLA: Learning to Navigate Socially Through Proactive Risk Perception - Technical Report for IROS 2025 RoboSense Challenge Social Navigation Track

### v1

@@ 1 Introduction
Social Navigation (SocialNav) denotes the capability of autonomous robots to navigate in human-shared environments while adhering to social conventions and maintaining culturally appropriate interpersonal distances [35]. Unlike conventional navigation paradigms that depend on pre-constructed static maps, SocialNav necessitates continuous adaptation to dynamic human behaviors while ensuring both collision-free trajectories and socially acceptable conduct. SocialNav Task is illustrated in Figure 1.
@@
Figure 1: SocialNav task illustration, adapted from [11]. In (a), the robot navigates toward a goal (blue dashed lines) while predicting human trajectories (red dashed line) and avoiding them. The robot uses depth input as shown in (b). (c) offers a top-down map for reference, which is not used by the robot.
@@
Contemporary approaches to this challenge predominantly employ reinforcement learning (RL) frameworks [25, 49, 22]. However, conventional RL-based methods frequently exhibit myopic behavior, focusing on immediate obstacle avoidance rather than anticipating future scenarios [5, 27]. While hierarchical architectures combining global path planners with local RL policies have shown promise [39], they inherently depend on comprehensive environmental knowledge [4], rendering them impractical for real-world deployment where such information is unavailable. The integration of human trajectory forecasting has demonstrated substantial benefits for dynamic collision avoidance [28, 38], though primarily in outdoor autonomous driving contexts [30, 29]. Indoor environments present distinct challenges, including constrained maneuvering space and elevated collision probabilities [39]. The recently proposed Falcon framework addresses these limitations through future-aware mechanisms, incorporating a Social Cognition Penalty and Spatial-Temporal Precognition Module to enhance trajectory prediction capabilities.
@@
In the IROS 2025 RoboSense Challenge Social Navigation Track, participants are tasked with developing navigation policies based exclusively on egocentric RGBD observations and odometry data. The benchmark dataset, Social-HM3D [11], provides photo-realistic large-scale indoor scenes populated with collision-aware human agents, offering a rigorous testbed for evaluating social navigation algorithms. Building upon the Falcon [11], we introduce a Proactive Risk Perception Module that enhances the agent’s capability to assess and respond to collision risks in human-populated environments, as illustrated in Figure 2. Our module employs a neural network to predict continuous risk scores for each nearby human based on their relative distances, with graduated risk levels spanning from safe zones to immediate danger regions. By incorporating distance-weighted loss functions that emphasize high-risk scenarios, our approach enables more nuanced spatial reasoning and anticipatory collision avoidance strategies.
@@ 2 Related Work
Social Navigation. The SocialNav paradigm [39, 10] extends traditional point-goal navigation by incorporating dynamic human agents into navigation scenarios, first formalized in the iGibson SocialNav Challenge [52]. Early implementations featured simplified human representations with constrained motion patterns. Contemporary research benefits from advanced simulation platforms such as Habitat 3.0 [40], which provides photorealistic rendering and naturalistic human animations. The SocialNav domain has attracted extensive research attention across robotics, computer vision, and behavioral studies [44, 37, 56, 57, 51, 54, 51, 55]. Foundational work in multi-agent collision avoidance [46, 45, 32] and dynamic obstacle navigation [3] has evolved to specifically address human-robot interaction challenges [13, 9, 7]. Several approaches model interpersonal dynamics through spatio-temporal graph representations [33, 6] to capture evolving agent relationships. Recent investigations emphasize egocentric navigation in realistic settings [42, 34, 48]. Our work distinguishes itself by incorporating Proactive Risk Perception that complements trajectory prediction capabilities for enhanced collision avoidance.
@@
Auxiliary Learning in Navigation. Auxiliary task learning has proven effective for improving task performance in navigation [36]. Diverse auxiliary objectives have been explored [24, 31], including environmental property prediction and agent state forecasting [1, 24]. Recent work has leveraged privileged information such as human-agent spatial relationships as auxiliary signals [4].
@@
Trajectory Prediction. Human trajectory forecasting plays a fundamental role in autonomous system safety [43, 23, 15, 16, 15, 17, 14, 19, 20, 18]. Classical methodologies include physics-based models such as Social Force [9], which simulate social dynamics through force interactions. Contemporary approaches span three paradigms: physics-based methods deriving from Newtonian mechanics [8, 53], data-driven techniques learning from historical patterns [2], and intent-reasoning methods that model goal-directed behavior [47].
3 Method
@@
Figure 2: Method overview, adapted from [11]. Building upon the Falcon framework, our approach integrates a Proactive Risk Perception Module that operates alongside the main policy network. The policy network processes depth and GPS+Compass observations, guided by social cognition penalties. During training, the state encoder outputs are fed to both Falcon’s original Spatial-Temporal Precognition Module and our proactive risk perception module. The risk module predicts distance-based collision risks for each nearby human, generating an auxiliary loss that enhances the agent’s spatial awareness and collision avoidance capabilities.
@@ 3.1 Task Definition
We formulate the social navigation challenge as follows: an autonomous agent $a$ operates within an environment containing $N$ dynamic human entities, indexed as $i\in\{1,\dots,N\}$. Beginning from initial state $q_{a}\in Q$, the agent must generate a trajectory $\tau_{a}$ toward target state $g_{a}\in Q$ while maintaining collision-free paths with both static structures and moving humans. This objective can be expressed as:
@@ 3.2 Falcon Framework Overview
Our method builds upon the Falcon architecture [11], which comprises a main policy network trained with social cognition penalties and a spatial-temporal precognition module for auxiliary learning. We briefly review these components before introducing our contribution.
@@ 3.2.1 Main Policy Architecture
The policy network accepts depth imagery and relative goal coordinates as input at each timestep, outputting navigation actions. Visual features are extracted via a ResNet-50 [21] encoder, while goal information passes through a linear transformation. A 2-layer LSTM [12] processes temporal dependencies, with outputs feeding into actor and critic heads for action generation and value estimation respectively.
@@ 3.2.2 Auxiliary Task Learning
$\mathcal{L}_{\text{traj}}=\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\|\hat{\mathbf{P}}_{i}^{\hskip 1.0ptt+1:t+H}-\mathbf{P}_{i}^{\hskip 1.0ptt+1:t+H}\|^{2}.$ (9)
@@ 3.3.2 Architecture Design
$\phi_{\text{risk}}(\delta_{R})=\sigma(W_{2}\cdot\text{ReLU}(W_{1}\cdot\delta_{R})).$ (10)
@@ 3.3.3 Risk Score Formulation
$r_{i}^{\text{true}}=\begin{cases}1.0&\text{if }d_{i}<d_{\text{danger}},\\ \frac{d_{\text{safe}}-d_{i}}{d_{\text{safe}}-d_{\text{danger}}}&\text{if }d_{\text{danger}}\leq d_{i}<d_{\text{safe}},\\ 0.0&\text{if }d_{i}\geq d_{\text{safe}}.\end{cases}$ (11)
@@ 3.3.4 Integration with Falcon
$\mathcal{L}_{\text{total}}=\beta_{\text{main}}\mathcal{L}_{\text{main}}+\beta_{\text{aux}}(\mathcal{L}_{\text{count}}+\mathcal{L}_{\text{pos}}+\mathcal{L}_{\text{traj}})+\beta_{\text{risk}}\mathcal{L}_{\text{risk}}.$ (12)
@@ 4.1 Dataset
The IROS 2025 RoboSense Challenge Social Navigation Track employs the Social-HM3D benchmark [11], a large-scale photo-realistic dataset specifically designed for evaluating socially-aware navigation in dynamic indoor environments. This benchmark addresses critical limitations in existing datasets by providing diverse scene types with balanced human density and naturalistic agent behaviors.
Dataset Characteristics. Social-HM3D is constructed from 3D-reconstructed real-world indoor scenes sourced from the HM3D dataset [41], featuring 844 unique environments spanning residential, commercial, and public spaces. The dataset incorporates collision-aware human agents with goal-directed behaviors, enabling realistic evaluation of human-robot interaction dynamics. Key design principles include:
Goal-Oriented Trajectories: Human agents navigate between designated waypoints within each scene, creating natural movement patterns that vary in complexity based on environmental layout. This design contrasts with prior datasets that employ random walk behaviors [26, 48], providing more authentic interaction scenarios.
Naturalistic Motion Dynamics: Walking speeds are randomized within 0.8-1.2 times the robot’s velocity to reflect human variability. Agents employ mutual collision avoidance using the ORCA algorithm [46], alternating between locomotion and stationary periods based on task completion. Realistic skeletal animations enhance visual fidelity and temporal coherence.
@@
Challenge Requirements and Restrictions. Participants train navigation policies on the social-hm3d/train split, which provides access to scene geometry, human trajectories, and RGBD observations. During development, teams can leverage the social-hm3d/val split approximately 1000 episodes for local validation, enabling iterative algorithm refinement without accessing the held-out test data. The final evaluation is conducted on a private test set comprising approximately 500 episodes from previously unseen scenes in social-hm3d/test. This held-out evaluation protocol rigorously assesses generalization capabilities without scene-specific overfitting.
@@ 4.3 Experiment Setups
Our implementation builds upon the Falcon framework with the integrated Proactive Risk Perception Module. The visual encoder employs ResNet-50 [21] architecture for depth processing, while a 2-layer LSTM handles temporal feature extraction. The risk perception module adds minimal computational overhead with two fully-connected layers operating on the shared hidden state representation.
We adopt the DD-PPO algorithm [50] for policy optimization. The model is initialized from pretrained PointNav weights provided by the Habitat platform [41] and fine-tuned on the social navigation task. Training is conducted on 4 NVIDIA A40 GPUs with 8 parallel environments for approximately 75 million steps.
@@ 5 Conclusion
In this report, we present our technical submission to the IROS 2025 RoboSense Challenge Social Navigation Track. Building upon the Falcon framework [11], we introduce a Proactive Risk Perception Module that enhances collision avoidance capabilities. Our comprehensive evaluation on the Social-HM3D benchmark demonstrates the effectiveness of risk-aware auxiliary learning for social navigation. The competition results validate our approach, achieving a competitive 2nd place among 16 participating teams.

### v2

# Learning to Navigate Socially Through Proactive Risk Perception

@@ 1 Introduction
Social Navigation (SocialNav) denotes the capability of autonomous robots to navigate in human-shared environments while adhering to social conventions and maintaining culturally appropriate interpersonal distances [mavrogiannis2023core]. Unlike conventional navigation paradigms that depend on pre-constructed static maps, SocialNav necessitates continuous adaptation to dynamic human behaviors while ensuring both collision-free trajectories and socially acceptable conduct. SocialNav Task is illustrated in Figure 1.
@@
Figure 1: SocialNav task illustration, adapted from [gong2025cognition]. In (a), the robot navigates toward a goal (blue dashed lines) while predicting human trajectories (red dashed line) and avoiding them. The robot uses depth input as shown in (b). (c) offers a top-down map for reference, which is not used by the robot.
@@
Contemporary approaches to this challenge predominantly employ reinforcement learning (RL) frameworks [kapoor2023socnavgym, wang2024multi, hirose2024selfi]. However, conventional RL-based methods frequently exhibit myopic behavior, focusing on immediate obstacle avoidance rather than anticipating future scenarios [chen2019crowd, li2019sarl]. While hierarchical architectures combining global path planners with local RL policies have shown promise [perez2021robot], they inherently depend on comprehensive environmental knowledge [cancelli2023exploiting], rendering them impractical for real-world deployment where such information is unavailable. The integration of human trajectory forecasting has demonstrated substantial benefits for dynamic collision avoidance [liang2019peeking, nishimura2020risk], though primarily in outdoor autonomous driving contexts [liang2020garden, liang2020simaug]. Indoor environments present distinct challenges, including constrained maneuvering space and elevated collision probabilities [perez2021robot]. The recently proposed Falcon framework addresses these limitations through future-aware mechanisms, incorporating a Social Cognition Penalty and Spatial-Temporal Precognition Module to enhance trajectory prediction capabilities.
@@
In the IROS 2025 RoboSense Challenge Social Navigation Track, participants are tasked with developing navigation policies based exclusively on egocentric RGBD observations and odometry data. The benchmark dataset, Social-HM3D [gong2025cognition], provides photo-realistic large-scale indoor scenes populated with collision-aware human agents, offering a rigorous testbed for evaluating social navigation algorithms. Building upon the Falcon [gong2025cognition], we introduce a Proactive Risk Perception Module that enhances the agent’s capability to assess and respond to collision risks in human-populated environments, as illustrated in Figure 2. Our module employs a neural network to predict continuous risk scores for each nearby human based on their relative distances, with graduated risk levels spanning from safe zones to immediate danger regions. By incorporating distance-weighted loss functions that emphasize high-risk scenarios, our approach enables more nuanced spatial reasoning and anticipatory collision avoidance strategies.
@@ 2 Related Work
Social Navigation. The SocialNav paradigm [perez2021robot, francis2023principles] extends traditional point-goal navigation by incorporating dynamic human agents into navigation scenarios, first formalized in the iGibson SocialNav Challenge [xia2020interactive]. Early implementations featured simplified human representations with constrained motion patterns. Contemporary research benefits from advanced simulation platforms such as Habitat 3.0 [puig2023habitat], which provides photorealistic rendering and naturalistic human animations. The SocialNav domain has attracted extensive research attention across robotics, computer vision, and behavioral studies [singamaneni2024survey, moller2021survey, zhang2025mapnav, zhang2025multi, wu2025evaluating, zhang2024trihelper, wu2025evaluating, zhang2025nava]. Foundational work in multi-agent collision avoidance [van2011reciprocal, rvo, learned] and dynamic obstacle navigation [mpdyn] has evolved to specifically address human-robot interaction challenges [guzzi2013human, socialforces, sa-cadrl, li2025seeground, li2025_3eed, xu2025lima]. Several approaches model interpersonal dynamics through spatio-temporal graph representations [social-graph, socialattention] to capture evolving agent relationships. Recent investigations emphasize egocentric navigation in realistic settings [rudenko2020thor, martin2021jrdb, vuong2023habicrowd, chen2023clip2Scene, chen2023towards]. Our work distinguishes itself by incorporating Proactive Risk Perception that complements trajectory prediction capabilities for enhanced collision avoidance.
@@
Auxiliary Learning in Navigation. Auxiliary task learning has proven effective for improving task performance in navigation [mirowski2016learning]. Diverse auxiliary objectives have been explored [jaderberg2016reinforcement, lin2019adaptive], including environmental property prediction and agent state forecasting [agrawal2015learning, jaderberg2016reinforcement]. Recent work has leveraged privileged information such as human-agent spatial relationships as auxiliary signals [cancelli2023exploiting].
@@
Trajectory Prediction. The forecasting of trajectories plays a fundamental role in autonomous system safety [rudenko2020human, huang2023multimodal, hao2021matters, hao2022listen, hao2021matters, hao2023dual, hao2023uncertainty, hao2024mbfusion, hao2025mapfusion, hao2024your, hao2025safemap, kong2025largead, kong2023robo3d, xie2023robobev, survey_3d_4d_world_models]. Classical methodologies include physics-based models such as Social Force [socialforces], which simulate social dynamics through force interactions. Contemporary approaches span three paradigms: physics-based methods deriving from Newtonian mechanics [elnagar2001prediction, zernetsch2016trajectory], data-driven techniques learning from historical patterns [alahi2016social], and intent-reasoning methods that model goal-directed behavior [vasquez2016novel].
3 Methodology
@@
Figure 2: Method overview, adapted from [gong2025cognition]. Building upon the Falcon framework, our approach integrates a Proactive Risk Perception Module that operates alongside the main policy network. The policy network processes depth and GPS+Compass observations, guided by social cognition penalties. During training, the state encoder outputs are fed to both Falcon’s original Spatial-Temporal Precognition Module and our proactive risk perception module. The risk module predicts distance-based collision risks for each nearby human, generating an auxiliary loss that enhances the agent’s spatial awareness and collision avoidance capabilities.
@@ 3.1 Task Definition
We formulate the social navigation challenge as follows: an autonomous agent $a$ operates within an environment containing $N$ dynamic human entities, indexed as $i\in\{1,\dots,N\}$. Beginning from the initial state $q_{a}\in Q$, the agent must generate a trajectory $\tau_{a}$ toward the target state $g_{a}\in Q$ while maintaining collision-free paths with both static structures and moving humans. This objective can be expressed as:
@@ 3.2 Falcon Framework Overview
Our method builds upon the Falcon architecture [gong2025cognition], which comprises a main policy network trained with social cognition penalties and a spatial-temporal precognition module for auxiliary learning. We briefly review these components before introducing our contribution.
@@ 3.2.1 Main Policy Architecture
The policy network accepts depth imagery and relative goal coordinates as input at each timestep, outputting navigation actions. Visual features are extracted via a ResNet-50 [resnet] encoder, while goal information passes through a linear transformation. A 2-layer LSTM [lstm] processes temporal dependencies, with outputs feeding into actor and critic heads for action generation and value estimation respectively.
@@ 3.2.2 Auxiliary Task Learning
$\mathcal{L}_{\text{traj}}=\frac{1}{|\mathcal{M}|}\sum_{i\in\mathcal{M}}\|\hat{\mathbf{P}}_{i}^{\hskip 1.0ptt+1:t+H}-\mathbf{P}_{i}^{\hskip 1.0ptt+1:t+H}\|^{2},$ (9)
@@ 3.3.2 Architecture Design
$\phi_{\text{risk}}(\delta_{R})=\sigma(W_{2}\cdot\text{ReLU}(W_{1}\cdot\delta_{R}))~,$ (10)
@@ 3.3.3 Risk Score Formulation
$r_{i}^{\text{true}}=\begin{cases}1.0&\text{if }d_{i}<d_{\text{danger}},\\ \frac{d_{\text{safe}}-d_{i}}{d_{\text{safe}}-d_{\text{danger}}}&\text{if }d_{\text{danger}}\leq d_{i}<d_{\text{safe}},\\ 0.0&\text{if }d_{i}\geq d_{\text{safe}},\end{cases}$ (11)
@@ 3.3.4 Integration with Falcon
$\mathcal{L}_{\text{total}}=\beta_{\text{main}}\mathcal{L}_{\text{main}}+\beta_{\text{aux}}(\mathcal{L}_{\text{count}}+\mathcal{L}_{\text{pos}}+\mathcal{L}_{\text{traj}})+\beta_{\text{risk}}\mathcal{L}_{\text{risk}},$ (12)
@@ 4 Experiments
In this section, we first introduce the dataset and evaluation metrics used in the IROS 2025 RoboSense Challenge Social Navigation Track. We then present our experimental setups and report the competition results, comparing our method with other participating teams’ methods.
@@ 4.1 Dataset
We use the official data provided by the RoboSense Challenge 2025 [robosense_challenge_2025] held at IROS 2025. This competition builds upon the legacy of the RoboDepth Challenge 2023 [robodepth_challenge_2023, kong2023robodepth] at ICRA 2023 and the RoboDrive Challenge 2024 [robodrive_challenge_2024, xie2025robobev] at ICRA 2024, continuing the collective effort to advance robust and scalable robot perception. Each track in this competition is grounded on an established benchmark designed for evaluating real-world robustness and generalization [xie2025drivebench, gong2025cognition, li2024place3d, chu2024geotext-1652, liang2025pi3det].
Specifically, this task is built upon the Falcon benchmark [gong2025cognition] in Track 2, which assesses socially compliant navigation policies in dynamic human environments using realistic RGB-D simulations and human-robot interaction metrics. This task employs the Social-HM3D benchmark [gong2025cognition], a large-scale photo-realistic dataset specifically designed for evaluating socially-aware navigation in dynamic indoor environments. This benchmark addresses critical limitations in existing datasets by providing diverse scene types with balanced human density and naturalistic agent behaviors.
4.1.1 Dataset Characteristics
Social-HM3D is constructed from 3D-reconstructed real-world indoor scenes sourced from the HM3D dataset [ramakrishnan2021habitat], featuring 844 unique environments spanning residential, commercial, and public spaces. The dataset incorporates collision-aware human agents with goal-directed behaviors, enabling realistic evaluation of human-robot interaction dynamics. Key design principles include:
Goal-Oriented Trajectories: Human agents navigate between designated waypoints within each scene, creating natural movement patterns that vary in complexity based on environmental layout. This design contrasts with prior datasets that employ random walk behaviors [igibsonchallenge2021, vuong2023habicrowd], providing more authentic interaction scenarios.
Naturalistic Motion Dynamics: Walking speeds are randomized within 0.8-1.2 times the robot’s velocity to reflect human variability. Agents employ mutual collision avoidance using the ORCA algorithm [van2011reciprocal], alternating between locomotion and stationary periods based on task completion. Realistic skeletal animations enhance visual fidelity and temporal coherence.
@@ 4.1.1 Dataset Characteristics
4.1.2 Challenge Requirements & Restrictions
Participants train navigation policies on the social-hm3d/train split, which provides access to scene geometry, human trajectories, and RGBD observations. During development, teams can leverage the social-hm3d/val split approximately 1000 episodes for local validation, enabling iterative algorithm refinement without accessing the held-out test data. The final evaluation is conducted on a private test set comprising approximately 500 episodes from previously unseen scenes in social-hm3d/test. This held-out evaluation protocol rigorously assesses generalization capabilities without scene-specific overfitting.
@@ 4.3 Experiment Setups
Our implementation builds upon the Falcon framework with the integrated Proactive Risk Perception Module. The visual encoder employs ResNet-50 [resnet] architecture for depth processing, while a 2-layer LSTM handles temporal feature extraction. The risk perception module adds minimal computational overhead with two fully-connected layers operating on the shared hidden state representation.
We adopt the DD-PPO algorithm [wijmans2019dd] for policy optimization. The model is initialized from pretrained PointNav weights provided by the Habitat platform [ramakrishnan2021habitat] and fine-tuned on the social navigation task. Training is conducted on 4 NVIDIA A40 GPUs with 8 parallel environments for approximately 75 million steps.
@@ 5 Conclusion
In this report, we present our technical submission to the IROS 2025 RoboSense Challenge Social Navigation Track. Building upon the Falcon framework [gong2025cognition], we introduce a Proactive Risk Perception Module that enhances collision avoidance capabilities. Our comprehensive evaluation on the Social-HM3D benchmark demonstrates the effectiveness of risk-aware auxiliary learning for social navigation. The competition results validate our approach, achieving a competitive 2nd place among 16 participating teams.